# Uncertainty-Calibrated Test-Time Model Adaptation without Forgetting

### v1

@@ 5 Experiments
[…]
ViT(LN) 52.5 5.0 30,000 0
[…]
Datasets and Models. We conduct experiments on three benchmarks datasets for OOD generalization, i.e., ImageNet-C [9] (contains corrupted images in 15 types of 4 main categories and each type has 5 severity levels) and ImageNet-R [65] for image classification, and ACDC [66] for semantic segmentation. We use ResNet-50 (R-50) [1] and VitBase (ViT) [67] for ImageNet experiments, and Segformer-B5 [68] for ACDC [66] experiments. The models are trained on ImageNet or CityScapes [69] training set with stochastic depth regularization [63] and then tested on clean or the above OOD test sets.
Compared Methods. We compare with the following state-of-the-art methods. BN adaptation [32] updates batch normalization statistics on test samples. Tent [12] minimizes the entropy of test samples during testing. MEMO [14] maximizes the prediction consistency of different augmented copies regarding a given test sample. CoTTA [64] minimizes the cross entropy between the predictions of the student network and its mean teacher at test time. SAR [18] selects reliable samples to perform test time sharpness-aware entropy minimization. We denote EATA without weight regularization in Eqn. (9) as efficient test-time adaptation (ETA). More ablative methods can be found in Table VI.
Round 1 4 7 10 All
BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0${}_{(-4.7)}$
Tent (lifelong) 69.0 40.2 60.0 57.3 66.6 36.6 58.9 54.2 64.6 33.4 55.9 51.6 62.5 30.4 52.6 48.7 52.7${}_{(-4.0)}$
CoTTA 70.9 41.1 62.4 59.7 70.8 40.6 62.6 59.7 70.8 40.5 62.6 59.7 70.8 40.5 62.6 59.7 58.4${}_{(+1.7)}$
EATA 69.1 40.5 59.8 58.1 69.3 41.8 60.1 58.6 68.8 42.5 59.4 57.9 67.9 42.8 57.7 56.3 57.0${}_{(+0.3)}$
EATA-C 71.0 44.3 63.1 61.1 72.0 47.3 64.9 63.8 71.8 48.2 64.2 64.2 72.0 48.7 64.3 64.1 61.6${}_{(+4.9)}$
Implementation Details. For test time adaptation, we use SGD as the update rule, with a momentum of 0.9 and a batch size of 64. In EATA and ETA, the learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase on ImageNet, and 7.5×$\times 10^{-5}$ on ACDC, respectively (following Tent, SAR and CoTTA). In EATA-C, the learning rate is set to 0.005/0.1 for ResNet-50/ViTBase on ImageNet, and 0.0005 on ACDC, respectively. The sub-network is obtained via stochastic depth regularization [63] with a drop ratio of 0.2/0.6 for ImageNet/ACDC. For both EATA and EATA-C, we use 2,000/20 samples for ImageNet/ACDC to calculate $\omega(\theta_{i})$ in Eqn. (11). More details are put in Supplementary.
Results on ImageNet-C. From the results, our EATA and EATA-C significantly outperform the considered methods over both ResNet-50 and ViTBase regarding the classification accuracy, suggesting our effectiveness. With our sample-adaptive entropy loss, EATA achieves a large performance gain over the counterpart Tent (e.g., $33.4\%\rightarrow 50.5\%$ on Gaussian noise with ViTBase), verifying that removing samples with unreliable gradients and tackling samples differently benefits the test-time adaptation. More critically, EATA-C outperforms CoTTA and SAR consistently in all 15 corruption types (while EATA fails to achieve this), demonstrating the effectiveness to reduce model uncertainty with our consistency loss. Note that besides achieving strong OOD performance, EATA also alleviates the forgetting on ID samples (see Figure 3), showing the effectiveness of our anti-forgetting regularization without limiting the learning ability during adaptation (see also Table VI for ablation).
As for efficiency, the required average backward number of our EATA is 29,721 on ResNet-50, which is much fewer than the method that needs multiple data augmentations (i.e., MEMO is 50,000$\times$64) or that requires multiple rounds of optimizations (e.g., SAR is 68,145 on ResNet-50). Compared with Tent, EATA reduces the average #backward (e.g., 50,000$\rightarrow$29,721 on ResNet-50) by excluding samples with high prediction entropy and samples that are similar out of test-time optimization. In this sense, our method only needs to adapt for partial samples, resulting in higher efficiency. While EATA-C requires more forward passes, its forward passes on all test samples using the full network are gradient-free, and the more lightweight sub-network’s forward/backward passes are conducted only on the selected reliable and non-redundant samples, maintaining comparable efficiency. Note that although optimization-free methods (such as BN adaptation) do not need backward updates, their OOD performances are limited.
In terms of calibration, existing methods consistently exhibit severe overconfidence (e.g., MEMO and CoTTA are 35.63% and 27.59% on ECE with ViTBase), suggesting that overconfidence is a prevalent issue in the unsupervised test-time adaptation. Note that by filtering unreliable samples to mitigate error accumulation, EATA achieves better calibration compared with Tent (e.g., $11.7\%\rightarrow 10.5\%$ on Gaussian noise with ResNet-50). However, the miscalibration is yet significant. By further differentiating the reducible model uncertainty and the inherent data uncertainty in adaptation, our EATA-C is able to reduce the calibration error by relatively 59.8% on ResNet-50 and 64.9% on ViTBase compared with EATA, demonstrating the promising calibration effect of EATA-C on different datasets and various architectures. It is worth mentioning that while EATA-C enhances performance and efficiency, EATA-C further achieves consistent and substantial improvement over accuracy, efficiency, and calibration error compared with state-of-the-art methods, suggesting our superiority.
Results on ImageNet-R. From Table III, our EATA consistently achieves better performance than compared methods on both ResNet-50 and ViTBase, while significantly reducing the required backward propagation (e.g., 30,000normal-$\rightarrow$5,417 on ResNet-50 over Tent), demonstrating the effectiveness of adapting on partial samples during testing. Our EATA-C further improves performance by a large margin (+6.0% over EATA on ViTBase), while achieving the best calibration. More critically, EATA-C is the only method that reduces calibration error on both ResNet-50 and ViTBase, suggesting the superiority in managing model uncertainty and data uncertainty differently for calibrated test-time adaptation.
Results on CityScapes-to-ACDC. Following [64], we evaluate our method on semantic segmentation task in the lifelong adaptation scenario. From Table IV, EATA achieves comparative performance by filtering unreliable predictions and constraining the important weights from drastic change, while its counterpart, Tent, severely suffers from error accumulation. Moreover, by differentiating model uncertainty and data uncertainty in test-time adaptation, EATA-C achieves state-of-the-art performance, outperforming EATA by 4.6% and CoTTA by 3.2% in mIoU. More importantly, EATA-C can progressively accumulate knowledge over adaptation to improve its performance. In the first and tenth rounds of adaptation, EATA-C achieves 59.8% vs. 62.3% in average mIoU on the four adverse datasets, further verifying the effectiveness of our method.
5.2 Demonstration of Preventing Forgetting
5.3 Ablation Studies
Effect of Components in $S({\bf x})$ (Eqn. 8). Our EATA accelerates test-time adaptation by excluding two types of samples out of optimization: 1) samples with high prediction entropy values (Eqn. 3) and 2) samples that are similar (Eqn. 8). We ablate both of them in Table V. Compared with the baseline $S({\bf x})\small{=}1$ (the same as Tent), introducing $S^{ent}({\bf x})$ in Eqn. (3) achieves better accuracy and fewer backwards (e.g., 49.6% (37,636) vs. 33.4% (50,000) on level 5). This verifies our motivation in Figure 2 that some high-entropy samples may hurt the performance since their gradients are unreliable. When further removing some redundant samples that are similar (Eqn. 8), our EATA further reduces the number of back-propagation (e.g., 37,636$\rightarrow$28,168 on level 5) and achieves comparable OOD error (e.g., 50.4% vs. 49.6%), demonstrating the effectiveness of our sample-efficient optimization strategy.
Effect of Components in EATA-C. By further differentiating between model uncertainty and data uncertainty based on EATA, our EATA-C achieves a substantial improvement in performance, efficiency and calibration. We conduct ablation to verify their effectiveness in Table VI. Note that adapting with simply the consistency loss in Eqn. (12) already leads to a strong baseline. By re-calibrating confidence on different samples via Eqn. (14), we further mitigate the calibration error, e.g., 3.3% (Exp1$\rightarrow$2.5(Exp3) on ECE. Compared with Exp3, Exp6 maintains comparable calibration while improving accuracy (i.e., 65.9$\rightarrow$66.4%), demonstrating that our anti-forgetting regularizer does not limit OOD adaptation. When further introducing the sample selection in Eqn. (17), EATA-C is able to effectively reduce both forward passes (i.e., 50,000$\times$2$\rightarrow$82,364) and backward passes (50,000$\rightarrow$32,364) by a large margin. Note that although accuracy and calibration are slightly compromised in EATA-C, due to the robustness of our consistency loss on corrupted images compared with Tent (see Figure 7), its improvement over the source model (65.9% vs.39.8% on accuracy, and 3.3% vs.7.5% on ECE) is still significant.
Figure 6: Effect of different stochastic depth ratios in EATA-C. Results obtained on ImageNet-C(Gaussian, level 5) with ViTBase.
Method Level 3 Level 5
Acc. (%) #Backwards Acc. (%) #Backwards
Baseline ($S({\bf x})\small{=}1$) 68.8 50,000 33.4 50,000
  $+S^{ent}{({\bf x})}$ (Eqn. 3) 70.7 45,302 49.6 37,636
  $+S{({\bf x})}$ (Eqn. 8) 70.8 36,057 50.4 28,168
Entropy Constant $E_{0}$ in Eqn. (3). We evaluate our EATA with different $E_{0}$, selected from {0.20, 0.25, 0.30, 0.35, 0.40, 0.45, 0.50, 0.55}×$\times\ln 10^{3}$, where $10^{3}$ is the class number of ImageNet. From Figure 6, our EATA achieves excellent performance when $E_{0}$ belongs to $[0.4,0.5]$. Either a smaller or larger $E_{0}$ would hamper the performance. The reasons are mainly as follows. When $E_{0}$ is small, EATA removes too many samples during adaptation and thus is unable to learn enough knowledge from the remaining samples. When $E_{0}$ is too large, some high-entropy samples would take part but contribute unreliable and harmful gradients, resulting in performance degradation. As larger $E_{0}$ leads to more backward passes, we set $E_{0}$ to 0.4×$\times\ln 10^{3}$ for efficiency-performance trade-off and fix the proportion of 0.4 for all other ImageNet experiments.
TABLE VI: Effects of components in EATA-C. Results obtained on 15 corruptions of ImageNet-C (severity level 5) with VitBase in single-domain adaptation scenario, i.e., the model parameters are reset before adapting to a new corruption. CL denotes consistency loss. ER denotes min-max entropy regularization. FR denotes anti-forgetting Fisher regularization. SS denotes active sample selection.
Experiment CL ER FR SS Average
Acc. ECE #Forwards #Backwards
1 65.6 3.3 50,000$\times$2 50,000
2 53.4 4.1 50,000$\times$2 50,000
3 65.9 2.5 50,000$\times$2 50,000
4 66.0 3.3 50,000$\times$2 50,000
5 53.0 4.5 50,000$\times$2 50,000
6 66.4 2.6 50,000$\times$2 50,000
7 65.6 4.4 83,182 33,182
8 52.9 3.4 73,323 23,323
9 (ETA-C) 65.3 3.4 82,015 32,015
10 66.2 4.3 83,429 33,429
11 53.2 3.4 73,723 23,723
12 (EATA-C) 65.9 3.3 82,364 32,364
@@ 5.4 More Discussions
Results on Single Sample Adaptation (Batch Size $B=1$). In our proposed methods, we conduct test-time learning over a batch of test samples each time. When batch size $B$ equals 1, directly applying Tent and our ETA on models with batch normalization layers may fail (see Table VII), but it is not a very significant issue in practice. Actually, one can address this by either maintaining a sliding window $\{{\bf x}_{i}\}_{i=t-L}^{t}$ including $L$ previous samples and the current test sample ${\bf x}_{t}$, or data augmentation techniques adopted in existing single sample TTA methods ([14, 33]). The first way is simple to implement and incurs only minor extra costs. From Table VII, our ETA with sliding window (ETA-wnd) works well and consistently outperforms the Tent counterpart.
Wall-clock Time Speed-up of EATA and EATA-C. In the current PyTorch version, gradient computation is conducted on the full mini-batch, even if instance-wise masks are applied. To benefit from active sample selection, in EATA, we perform one forward-only pass with $B$ samples, and one extra forward and backward pass with $B^{\prime}$ selected samples. We achieve wall-time speed-up when setting $E_{0}$ in Eqn. (3) to $0.1\times\ln C$, where $C$ is the number of task classes. For ViTBase on ImageNet-C (Gaussian noise, level=5) with one Titan Xp GPU, the actual run time is 518s for Tent (33.3% accuracy), 346s for EATA (39.5% accuracy), and 402s for EATA-C (51.3% accuracy). It’s also worth mentioning that in EATA-C, the sub-network’s computation is not accelerated even if we select partial samples for each layer due to a similar setback. Actually, this is a more engineering problem, and an ideal implementation should further speed up both EATA and EATA-C.
@@
Performance under Mixed-and-Shifted Distributions. We evaluate Tent and our EATA/EATA-C on mixed ImageNet-C (level=3 or 5) that consists of 15 different corruption types/distribution shifts (totaling 750k images). Results in Table VIII show the stability of EATA and EATA-C in large-scale and complex TTA scenarios.
TABLE VII: Effectiveness of ETA under sliding window strategy (with different window length $L$) for single sample adaptation with ResNet-50. We report corruption accuracy (%) on ImageNet-C (Gaussian noise, level 5).
Base ResNet-50 Tent ETA Tent-wnd ($L\small{=}32)$ ETA-wnd ($L\small{=}32$) Tent-wnd ($L\small{=}64$) ETA-wnd ($L\small{=}64$)
2.2 0.1 0.1 28.1 30.8 29.5 32.4
TABLE VIII: Comparison with Tent [12] w.r.t. corruption accuracy (%) with mixture of 15 corruption types on ImageNet-C with ViTBase.
 Severity Source Tent EATA (ours) EATA-C (ours)
Level=3 63.0 69.8${}_{(+6.8)}$ 71.1${}_{(+8.1)}$ 72.4${}_{(+9.4)}$
Level=5 39.8 47.0${}_{(+7.2)}$ 58.2${}_{(+18.4)}$ 60.4${}_{(+20.6)}$
@@
Figure 7: Comparisons of performance, calibration and stability using consistency loss in Eqn. (12) or entropy minimization loss (i.e., Tent) on ImageNet-C (Gaussian noise, severity level=5) with ViTBase. In Left, we split the dataset into 80% and 20% slices to conduct adaptation and to evaluate the adapted model. We report the performance and the calibration of the adapted model every 75 batches. In the Middle and Right, we evaluate the stability of using consistency loss and Tent under various combinations of learning rates and adaptation steps per batch. Consistency loss achieves substantially higher OOD performance and better stability while maintaining lower ECE.
Advantage of Consistency Loss over Tent [12]. We conduct a comprehensive comparison of the performance, calibration and stability between the use of consistency loss, as defined in Eqn. (12), and entropy minimization loss (i.e., Tent [12]) to reduce uncertainty at test time. The experiments are conducted on ImageNet-C (Gaussian noise, level 5) with ViTBase.
In Figure 7(a), we partition the dataset into 80% and 20% for adaptation and evaluation, respectively, to show how adaptation impacts performance and calibration over time. Analyzing Figure 7(a), we discern several key findings: 1) consistency loss consistently demonstrates superior performance and calibration throughout adaptation. 2) consistency loss is more sample-efficient, where adapting with as few as 75 batches can significantly outperform Tent [12] that adapts with 300 batches. 3) Tent [12] shows rapid degradation in performance and calibration after convergence, which indicates that it tends to over-fit. In contrast, consistency loss maintains stable performance and calibration after convergence and continuously exhibits strong generalization throughout adaptation.
In Figures 7(b) and 7(c), we study the stability of consistency loss and Tent [12] across various combinations of learning rates and adaptation steps following [71]. As illustrated in Figure 7(b), consistency loss demonstrates remarkable performance stability and benefits from the increment of adaptation steps (e.g., $55.8\%\rightarrow 58.6\%$, the best performance under 1 and 5 adaptation steps, respectively). In contrast, Tent [12] is sensitive to the combination of learning rate and adaptation steps, where its performance may deteriorate to as low as 1%. Such a phenomenon further highlights Tent [12]’s tendency to over-fit. In this sense, consistency loss emerges as the superior choice considering performance, calibration, stability and sample efficiency.

### v2

@@ 5.3 Ablation Studies
7 48.9 5.9 83,583 33,583
8 42.9 4.4 74,705 24,705
9 (ETA-C) 48.9 5.4 82,882 32,882
10 48.8 5.1 82,952 32,952
11 42.3 3.8 74,226 24,226
12 (EATA-C) 49.0 4.6 82,492 32,492
Similarity Threshold $\epsilon$ in Eqn. 7. We use $\epsilon$ to select diverse samples for TTA. From Figure 6, EATA maintains stable accuracy across a wide range of $\epsilon\in[0.01,0.08]$, showcasing insensitivity, while a smaller $\epsilon$ removes significantly more samples and improves computational efficiency. We set $\epsilon{=}0.05$ without careful tuning. More results on efficiency (i.e., time and memory usage) of EATA and EATA-C with varying $\epsilon$ are provided in Table VII.
@@
Factor $\alpha$ for Entropy Regularizer in Eqn. 16. We directly set $\alpha{=}0.1$ to align the magnitudes of consistency loss and entropy regularization loss for EATA-C without careful tuning. From Figure 9, increasing $\alpha$ within $[0,0.1]$ effectively reduces more ECE while maintaining stable accuracy, verifying its efficacy. However, when $\alpha$ exceeds 0.1, the entropy regularization loss dominates the adaptation, which leads to a gradual decline in accuracy.
Factor $\beta$ for Fisher Regularizer in Eqn. 16. From Figure 9, compared to ETA-C (i.e., setting $\beta{=}0$), introducing the fisher regularizer consistently achieves better accuracy. Moreover, once activated, the performance of EATA becomes largely insensitive to $\beta$ within the tested range of $[20,140]$, highlighting its robustness.
@@ 5.4 More Discussions
Efficiency Analysis of EATA and EATA-C. We evaluate the efficiency of our methods by including a more comprehensive comparison of time and memory usage for TTA, as in Table VII. The results reveal the following: 1) Adaptation Time: EATA and EATA-C require significantly less adaptation time than most baseline methods, while maintaining competitive or superior accuracy. For example, on ViT-Base, EATA-C improves accuracy from 48.8% (ROID) to 56.8% with a reduction in adaptation time from 181.6 seconds to 114.9 seconds. Moreover, by setting a stricter threshold $\epsilon$ to filter redundant test samples, EATA and EATA-C can be further accelerated while maintaining performance. For example, using $\epsilon{=}0.02$, EATA on ViT-Base increases accuracy from 33.4% (Tent) to 50.7% while reducing adaptation time from 106.1 seconds to 85.1 seconds; 2) Memory Usage: Our methods also demonstrate efficient memory utilization, where EATA and EATA-C consume substantially less memory compared to all competing TTA methods, e.g., on ResNet-50, memory usage decreases from 5417.7MB (SAR) to 2205.5MB (EATA-C, $\epsilon{=}0.02$) while accuracy increases from 29.6% to 35.9%. This memory reduction is achieved through our active sample selection strategy, which reduces the number of samples involved in backpropagation. Note that the current Pytorch implementation does not support instance-wise gradient computation, thus an ideal implementation should further speed up both EATA and EATA-C. See more discussions on our implementation details in Appendix A.
TABLE VII: Efficiency comparison regarding wall-clock time and peak memory usage on ImageNet-C (Gaussian, severity level 5) with an A100 GPU.
Method ResNet-50 ViT-Base
Acc. (%) Time (s) Mem. (MB) Acc. (%) Time (s) Mem. (MB)
Source 1.8 54.6 771.7 12.9 55.7 816.6
Tent [12] 28.0 99.9 5417.6 33.4 106.1 7433.2
SAR [18] 29.6 149.0 5417.7 43.1 167.3 7433.2
ROID [64] 36.7 190.2 9315.3 48.8 181.6 12321.4
CoTTA [16] 19.9 241.0 12196.7 45.6 229.9 22296.5
TEA [65] 18.5 2266.1 15942.0 46.9 - -
MEMO [14] 6.8 36329.2 8154.3 32.2 39918.8 11061.1
EATA ($\epsilon{=}0.02$) 35.2 93.0 2285.2 50.7 85.1 4234.6
EATA-C ($\epsilon{=}0.02$) 35.9 95.2 2205.5 55.0 88.8 4014.4
EATA (Ours) 35.6 106.5 3693.0 50.5 108.7 5887.6
EATA-C (Ours) 37.2 122.4 3358.7 56.8 114.9 5786.6
TABLE VIII: Reliability of data uncertainty indicator. We report sub-model Acc. (%) on ImageNet-C(Gaussian, level 5) after adapting to $B$ batches (batch size 64) with ViT-Base. “#Uncertain” are samples with disagreed predictions. “Indicator Acc.” is the ratio of these samples misclassified by sub-model.
Metric
Source
EATA-C
EATA-C
EATA-C
EATA-C
EATA-C
(B=150B{=}150) (B=300B{=}300) (B=450B{=}450) (B=600B{=}600) (B=750B{=}750)
Model Acc. (%) 9.3 47.0 50.2 50.7 51.9 52.0
#Uncertain 26144 15697 14682 14149 13648 13722
Indicator Acc. (%) 97.1 90.0 88.9 89.3 88.6 89.1
Figure 10: Calibration comparison of each confidence interval on ImageNet-C(Fog, level 5) with ViT-Base. Results are visualized following [44].
Effectiveness of Data Uncertainty Indicator. We evaluate the effectiveness of our data uncertainty indicator, i.e., Eqn. (15), throughout TTA. The results are detailed in Table VIII: 1) Consistently High Indicator Accuracy: Across various adaptation stages, our indicator continues to reliably identify uncertain samples, on which the sub-model’s prediction is likely to be incorrect. Specifically, the indicator maintains around 90% accuracy, suggesting its effectiveness even in the early stage of adaptation. This reliability allows us to apply entropy maximization on these uncertain data to improve calibration without hindering adaptation; 2) Reduced Uncertain Samples Over Time: The model’s initial poor performance mainly leads to a higher number of uncertain samples, including inherently difficult data for discrimination and data the model has yet to learn. However, model uncertainty quickly explains away during TTA (i.e., within one-fifth of the data stream), leading to a stabilized number of uncertain samples that reflects irreducible data uncertainty.
@@
Performance under Mixed-and-Shifted Distributions. We evaluate Tent and our EATA/EATA-C on mixed ImageNet-C (level=3 or 5) that consists of 15 different corruption types/distribution shifts (totaling 750k images). Results in Table IX show the stability of EATA and EATA-C in large-scale and complex TTA scenarios.
TABLE IX: Comparison with Tent [12] w.r.t. corruption accuracy (%) with mixture of 15 corruption types on ImageNet-C with ViT-Base.
 Severity
Source
Tent
EATA (ours)
EATA-C (ours)
Level=3 63.0 69.8(+6.8) 71.1(+8.1) 72.4(+9.4)
Level=5 39.8 47.0(+7.2) 58.2(+18.4) 60.4(+20.6)
@@
Figure 11: Comparisons of performance, calibration and stability using consistency loss in Eqn. (12) or entropy minimization loss (i.e., Tent) on ImageNet-C (Gaussian noise, severity level=5) with ViT-Base. In Left, we split the dataset into 80% and 20% slices to conduct adaptation and to evaluate the adapted model. We report the performance and the calibration of the adapted model every 75 batches with a batch size of 64. In the Middle and Right, we evaluate the stability of consistency loss and Tent under various combinations of learning rates and adaptation steps per batch. Consistency loss achieves substantially higher OOD performance and better stability while maintaining lower ECE.
Calibration Across Confidence Intervals. EATA-C aims to achieve a favorable balance among accuracy, calibration, and efficiency. In EATA-C, we discard high-entropy samples (termed active sample selection) mainly to improve computational efficiency. While some high-entropy samples might benefit from further calibration, they typically yield unreliable pseudo-labels, which can negatively impact the stability and effectiveness of TTA (see Table VI). Instead of directly calibrating on these samples, we show that focusing adaptation and calibration on only low-entropy samples can also improve the calibration on high-entropy ones, as in Figure 10, while improving TTA efficiency and effectiveness.
Advantage of Consistency Maximization over Tent [12]. We conduct a comprehensive comparison w.r.t. performance, calibration, and stability between the use of consistency loss, as defined in Eqn. (12), and Tent [12] to reduce uncertainty during testing. From Figure 11(a), we have the following observations: 1) Consistency loss consistently demonstrates superior performance and calibration throughout adaptation. 2) Consistency loss is more sample-efficient, where adapting with as few as 75 batches can significantly outperform Tent [12] that adapts with 300 batches. 3) Tent [12] shows rapid degradation in performance and calibration after convergence. In contrast, consistency loss maintains stable performance and calibration after convergence and continuously exhibits strong generalization throughout TTA.
We further evaluate the stability of consistency loss and Tent [12] across various combinations of learning rates and adaptation steps in Figures 11(b) and 11(c) following [73]. The results highlight that our consistency loss demonstrates remarkable stability and benefits from increased adaptation steps (e.g., $55.8\%\rightarrow 58.6\%$, the best performance under 1 and 5 adaptation steps, respectively). In contrast, Tent [12] is highly sensitive to the combination of learning rate and adaptation steps, where its performance may deteriorate to as low as 1%, further indicating its tendency to overfit. These findings collectively underscore the superiority of our consistency loss regarding performance, calibration, and stability, making it a more robust choice for TTA.